A curve has an equation given by x2 – 2xy + y2 = 6.

a) Obtain an expression for dx dy in terms of x and y.

b) What is the gradient of the curve at (2, 3)? 


1
Expert's answer
2021-09-14T11:26:51-0400

a)

"F(x,y)=x^2 – 2xy + y^2 -6\\
F'_x(x,y)= 2x-2y\\
F'_y(x,y)= -2x+2y\\
\frac{dy}{dx}=−\frac{F'_x(x,y)}{F'_y(x,y)}= -\frac{2x-2y}{-2x+2y}\\
\frac{dy}{dx}=\frac{2y-2x}{2x-2y}\\
\implies \frac{dx}{dy}=\frac{2x-2y}{2y-2x}\\"


b)

"m= \frac{dy}{dx}\\
m=\frac{2y-2x}{2x-2y}\\
At (2, 3)\\
m=\frac{2*3-2*2}{2*2-2*3}\\
m=\frac{6-4}{4-6}\\
m=\frac{2}{-2}\\
m=-1"